# A note on an inequality involving the sides and medians in a triangle

### v1

@@ 1 Introduction
He made numerous numerical experiments which confirmed the conjectured relation, however, he was unable to find a definitive proof [^1] We challenge the interested readers to test their ideas on possible proof before reading further. He had communicated his conjecture in personal conversations with various colleagues including Mladen Vassilev-Missana, who relayed this story to us. Perhaps because such “Olympiad type problems” are not of great interest to the professional mathematicians or due to to the fact that it was not generally known to puzzle solvers, not many people undertook the task to find a proof.

### v2

@@ 1 Introduction
He made numerous numerical experiments which confirmed the conjectured relation, however, he was unable to find a definitive proof [^1] We challenge the interested readers to test their ideas on possible proof before reading further. He had communicated his conjecture in personal conversations with various colleagues including Mladen Vassilev-Missana, who relayed this story to us. Perhaps because such “Olympiad type problems” are not of great interest to the professional mathematicians or due to the fact that it was not generally known to puzzle solvers, not many people undertook the task to find a proof.
@@ 4 Additional results and discussion
Remark 1.
It seems likely that inequalities (18) and (19) are necessary but not sufficient for the fulfillment of (1) and (2).
If we consider the splitters (the Cevians that pass through the Nagel point and bisect the perimeter of the triangle), for a triangle with sides $a\leq b\leq c,$ we have for their lengths:
$n_{a}$ $=\sqrt{s}\sqrt{\frac{(c-b)^{2}}{a}+s-a};$
$n_{b}$ $=\sqrt{s}\sqrt{\frac{(c-a)^{2}}{b}+s-b};$
$n_{c}$ $=\sqrt{s}\sqrt{\frac{(b-a)^{2}}{c}+s-c},$
where $s=\frac{a+b+c}{2}$ is the semiperimeter. It is not hard to show that $n_{a}-n_{b}\geq 0,$ and $n_{b}-n_{c}\geq 0.$ For the first consider the expression
$\frac{ab(n_{a}^{2}-n_{b}^{2})}{s}=(b(c-b)^{2}+ab(b-a)-a(c-a)^{2})=(b-a)(a+b-c)^{2}\geq 0.$
Hence, $n_{a}-n_{b}\geq 0$ since $a,b,n_{a}+n_{n}$ and $s$ are all positive. Analogously, $n_{b}-n_{c}\geq 0$ is equivalent to
$\frac{bc(n_{b}^{2}-n_{c}^{2})}{s}=(c(c-a)^{2}+bc(c-b)-b(b-a)^{2})=(c-b)(c+b-a)^{2}\geq 0.$
Therefore, (18) is fulfilled.
For (19), let us first establish that $B=b\,n_{b}\geq a\,n_{a}=A.$
We have to show that
$b(c-a)^{2}+b^{2}(s-b)\geq a(c-b)^{2}+a^{2}(s-a)$
This inequality, however is equivalent to
$(2c-a-b)(b-a)s\geq 0,$
which in view of $a\leq b\leq c$ is true. Thus, $B\geq A.$ For $B\geq C=c\,n_{c},$ we have to show
$b(c-a)^{2}+b^{2}(s-b)\geq c(b-a)^{2}+c^{2}(s-s),$
This in turn is equivalent to
$(c+b-2a)(c-b)s\geq 0.$
The last, in view of $a\leq b\leq c$ is obviously true.
Hence, (19) is also fulfilled.
Let us consider the obtuse triangle with sides $a=4,b=16,c=18.$ For it we have, $n_{a}=4\sqrt{19},n_{b}=\frac{1}{2}\sqrt{61}\sqrt{19},n_{c}=3\sqrt{19}.$ A quick calculation shows that $C>A$ and both (1) and (2) fail to hold in this case.